OpenStudy (anonymous):

what is the distance between (-5,2) and (-9,2)

OpenStudy (anonymous):

The distance between (-5,2) and (-9,2) can be found by using Pythagorean Theorem. just draw it on a graph and connect the points to make a line segment, then you make a right triangle with the line segment. Find the lengths of the sides that share the right angle then use A^2 + B^2 =C^2 then do the squared root of C^2 and you will have the distance between the coordinates.

OpenStudy (anonymous):

Wow. Thanks so much. Now I remember!

OpenStudy (anonymous):

Thats good to hear :D

OpenStudy (anonymous):

\[d=\sqrt{\left( x2-x1 \right)^{2}+\left( y2-y1 \right)^{2}}\]

OpenStudy (anonymous):

Graphically 

OpenStudy (anonymous):

Its very simple, consider (-5,2) as A and (-9,2) as B where \[x1=-5 , y1=2, x2=-9, y2=2\], now there is a distance formula for straight line i.e \[d=\sqrt{(x2-x1)^2+(y2-y1)^2}\] \[\sqrt{(-9-(-5))^2+(2-2)^2}\] \[\sqrt{(-4)^2+0^2}\] \[\sqrt{16}=4\]